Question: Assume that the reserve ratio in a country is 16%. Calculate the simple money multiplier?
1
Expert's answer
2018-01-10T14:47:07-0500
If the reserve ratio in a country is rr = 16%, then the simple money multiplier is m = 1/rr = 1/0.16 = 6.25.